A box of cereal is pushed with an applied force of 2.4 N [W] across a table. If the cereal box has a mass of 2.5 kg and the box accelerates at 0.41 m/s^2 [W], The coefficient of friction for the table top is___.(give your answers with the correct number of significant digits)

Note: Let West be positive


1
Expert's answer
2021-07-08T09:57:53-0400

According to Newton's second law:


"ma=F-f,\\
ma=F-\mu mg,\\
\mu=\frac{F-ma}{mg}=0.06."